OpenStudy (anonymous):

Write y = –2/5x – 3 in standard form using integers.

OpenStudy (anonymous):

2x + 5y = –15 2x + y = –3 2x + 5y = –3 x + y = –15

zepdrix (zepdrix):

\[\large y=-\frac{2}{5}x-3\]So the first thing we have to do is get rid of the fraction since they want all of the coefficients to be INTEGERS. :) Hmm it looks like the largest denominator is a 5. So let's multiply both sides by 5.\[\large 5(y)=\left(-\frac{2}{5}x-3\right)5\]We'll have to distribute the 5 to each term on the right side.

zepdrix (zepdrix):

\[\large 5y=-5\cdot \frac{2}{5}x-3\cdot5\]

zepdrix (zepdrix):

From here, the 5's will cancel out on the x term, just as we wanted. And the 3 x 5 will give us 15.\[\large 5y=-\cancel{5}\cdot \frac{2}{\cancel{5}}x-15\] Giving us,\[\large 5y=-2x-15\]

zepdrix (zepdrix):

Then to get it into standard form, I think all we need to do is ADD 2x to each side right? :)

OpenStudy (anonymous):

Ahh thank you for explaining! (: And yes

OpenStudy (anonymous):

@zepdrix so which answer is it? (:

zepdrix (zepdrix):

Adding 2x to each side gives us, \[\large 5y+2x=\cancel{-2x}-15\cancel{+2x}\]

OpenStudy (anonymous):

So it's A! (:

OpenStudy (anonymous):

Can I ask you a question or two more @zepdrix ? I like how you explain it

OpenStudy (anonymous):

The table shows the height of a tree as it grows. What equation in slope-intercept form gives the tree's height at any time? Time Height (months) (inches) 2 14 4 23 6 32 8 41 y = 9x + 2 y = 9/2x + 5 y = 9/2x y = 5x + 9/2

OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (–7, 9); m = 4/5 y + 7 = 4/5(x – 9) y – 7 = 4/5(x + 9) y – 9 = 4/5(x + 7) y + 9 = 4/5(x – 7)

zepdrix (zepdrix):

Bahhh I gotta go do some laundry D: Soz! You should close this thread an open a new one! C: So people can find your questions easier instead of scrolling way way down :D heh